OpenStudy (anonymous):

write an equation for the linear function satisfying f(-5)=-1 and f(2)=4, y=?x+?

OpenStudy (tyteen4a03):

You are actually given 2 points. When x = -5, y = -1. When x = 2, y = 4. Now build the equation with these information.

OpenStudy (anonymous):

Am I supposed to find the slope and then use the point-slope formula?

OpenStudy (tyteen4a03):

@anicuhh Yup.

OpenStudy (anonymous):

I got f(x)=5/7x + 18/7 from that. I hope I did it right.

OpenStudy (tyteen4a03):

@anicuhh Remember to use parentheses. 5/7x means \(\frac{5}{7x}\), not\(\frac{5x}{7}\).

OpenStudy (anonymous):

Alright, thanks. c: